OpenStudy (anonymous):

the second angle of a triangle is 4 times as large as the first. the third angle is 5 more than the sum of the other two angles. find the measure of the second angle

Miracrown (miracrown):

We assume that the three angles are x, y, and z. x+y+x=180 because they are interior angles of a triangle. y=4x since the second angle is 4 times the first one. z=5+x+y because the third one is 5 more than the sum of the other two.

Miracrown (miracrown):

x+y+z=180 y=4x z=5+x+y

Miracrown (miracrown):

I got y=70. So the second angle is 70 degrees.

OpenStudy (mathstudent55):

Call the first angle x. The second angle is 4 times the first angle, so the second angle is 4x. The third angle is 5 more than the sum of the other two angles, so the third angle is: x + 4x + 5 The sum of the three measures is 180, so you get: (x) + (4x) + (x + 4x + 5) = 180 Solve for x, which gives you the first angle. Then multiply x by 4 to get the second angle.